OpenStudy (anonymous):

Help please... A survey of high school juniors found that 82% of students plan on attending college. If you pick three students at random, what is the probability that at least two plan on attending college? Round to the nearest percent. 36% 91% 45% 9%

OpenStudy (anonymous):

you doin e2020 ?

OpenStudy (anonymous):

I guess it would be 91% if calculated right xD

OpenStudy (anonymous):

91.45 To be exactly right

OpenStudy (anonymous):

How do you figure that, @Camtron ?

OpenStudy (anonymous):

I don't know exactly how to explain it to you :/

OpenStudy (anonymous):

How'd you solve?

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

do you know any other answer?

OpenStudy (anonymous):

well there you go xD

OpenStudy (anonymous):

haha Thanks! That makes sense... haha

OpenStudy (anonymous):

>.> Where did the response go?

OpenStudy (anonymous):

P(at least 2) is P(2) + P(3) P(2) = (3C2)(1 - 0.82)(0.82)^2 = 0.363096 P(3) = (0.82)^3 = 0.551368 P(at least 2) = 0.363096 + 0.551368 = 0.914464

OpenStudy (anonymous):

So, rounded, that's 91%

OpenStudy (anonymous):

Ahhhhh OK! I got it. I see where I messed up. lol I put 36% as the answer.

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

uw! Good luck to you in all of your studies and thx for the recognition! @Peasant1